# Biologically Plausible Learning via Bidirectional Spike-Based Distillation

### v1

@@ 6 Conclusion
Human learning and cognition emerge from bidirectional processes that integrate bottom-up sensory perception with top-down memory recall. In this framework, the feedforward network supports perception and decision-making by transforming sensory stimuli into conceptual representations, while the feedback network facilitates memory recall by reconstructing stimuli from semantic concepts. Inspired by this principle, we propose a novel bidirectional learning framework in which the feedforward and feedback networks are jointly trained by distilling their hidden feature representations from one another. This approach leverages the properties of pyramidal neurons, which receive feedforward (perception) and feedback (learning) signals through their basal and apical dendrites, respectively. Extensive experiments across diverse network architectures and tasks show that the resulting learning algorithm achieves performance comparable to error backpropagation while yielding stronger adherence to biological fidelity. Specifically, BSD achieves competitive performance across tasks: within 3% of backpropagation on MNIST (99.44% vs 99.56%) and CIFAR-10 (84.13% vs 87.12%), and superior MSE on time-series forecasting tasks like Electricity (0.165 vs 0.175), demonstrating the effectiveness of our approach. The limitations and future directions are discussed in Appendix G.
@@
The authors have made extensive efforts to ensure the reproducibility of the empirical results reported in this paper. First, detailed dataset characteristics are documented in Appendix E.1. Second, comprehensive implementation details, covering network architectures, hyperparameter settings, training procedures, and evaluation metric explanations, are provided in Appendix E.2. Finally, the authors have included the source code for the proposed training algorithm as supplementary material and promised to release it publicly on GitHub upon acceptance.
@@ Appendix A Global Algorithm of BSD
28:   We calculate the loss of the last layer using cross-entropy: $\mathcal{L}_{\text{target}}\leftarrow\mathcal{L}_{\text{CE}}(\mathbf{v}_{L},\hat{\mathbf{v}}_{L})$
@@ Appendix B Additional Preliminaries
Spiking neurons can be categorized into two types based on their signal propagation mechanisms. One category, referred to as spike-based neurons, conveys information through spike trains. Discrete Spiking Neural Networks (SNNs) integrate spike inputs at each time step and generate a binary spike output when the integrated value exceeds a threshold. Continuous analog input signals are typically encoded (e.g., rate encoding, time encoding, or $\Delta$-encoding) into spike sequences. Representative models in this category include IF, PLIF(Fang et al., 2021), and LIF(Maass, 1997). Discrete SNNs are compatible with existing digital design processes, allowing the grouping and virtualization of large numbers of neurons on computational cores to achieve parallel architectures. These systems exhibit high scalability and efficient trainability, benefiting from sparse, event-driven computation with low power consumption during the inference phase. However, the discrete nature of spike signals limits their information capacity relative to traditional neuron models, which transmit continuous floating-point values directly(Pfeiffer & Pfeil, 2018).
The second type of spiking neuron architecture uses similar temporal equations as spiking neurons but propagates analog signals, referred to as analog-based neurons, such as LIAF(Wu et al., 2021). These signals can encode more information and exhibit lower response latency in certain tasks. However, hardware implementations of such neurons require analog circuits and because computations cannot be skipped during inference, the system exhibits higher power consumption (Hazan & Ezra Tsur, 2021).
@@ E.2.2 BSD-CNNs
Similarly to the MLP case, the feedback path, composed of Type 2 neurons, mirrors the structure of the feedforward path composed of Type 1 neurons. In the feedback path, we use upsampling as the reverse operation of the convolutional layers used in the feedforward path, ensuring that the number of Type 1 and Type 2 neurons in feedforward and feedback paths remains equal across all layers.
@@ F.3 Ablation study on autoencoders
F.4 Dynamics Of Synaptic Weight Alignment
@@
In this section, we examine the alignment dynamics between the feedforward synaptic weights ($\mathbf{W}$) and the backward synaptic weights ($\bm{\Theta}$) within a convolutional neural network trained using the BSD algorithm. The network was tasked with image classification on the CIFAR-10 dataset. We measured the cosine similarity between the weights connected to the basal dendrites of Type 1 and Type 2 neurons to quantify their alignment in each layer. In Figure 7, ’enc1’ through ’enc5’ denote the five feedforward convolutional layers and their corresponding backward upsampling layers, progressing from the shallowest to the deepest. The ’output’ label refers to the final fully connected layer. Figure 7 reveals that the weight alignment for the final fully connected layer undergoes a rapid initial increase, peaking within the first few epochs before stabilizing at a high cosine similarity value of approximately 0.7. In contrast, the weights of the convolutional and upsampling layers (’enc1’ to ’enc5’) consistently exhibit negligible alignment, with their cosine similarity values remaining close to zero throughout the entire training process. Such alignment behavior demonstrates that the feedforward and backward pathways do not converge toward symmetric weights under BSD. Thereby we confirm that the BSD algorithm adheres to criterion C1: Asymmetric synaptic weights for feedforward and feedback pathways.
F.5 t-SNE Visualization of Learned Feature Representations in BSD-trained CNNs
@@
Appendix G Limitations and Future Directions
G.1 Limitations

### v2

@@ 6 Conclusion
Human learning and cognition emerge from bidirectional processes that integrate bottom-up sensory perception with top-down memory recall. In this framework, the feedforward network supports perception and decision-making by transforming sensory stimuli into conceptual representations, while the feedback network facilitates memory recall by reconstructing stimuli from semantic concepts. Inspired by this principle, we propose a novel bidirectional learning framework in which the feedforward and feedback networks are jointly trained by distilling their hidden feature representations from one another. This approach leverages the properties of pyramidal neurons, which receive feedforward (perception) and feedback (learning) signals through their basal and apical dendrites, respectively. Extensive experiments across diverse network architectures and tasks show that the resulting learning algorithm achieves performance comparable to error backpropagation while yielding stronger adherence to biological fidelity. Specifically, BSD achieves competitive performance across tasks: within 3% of backpropagation on MNIST (99.44% vs 99.56%) and CIFAR-10 (84.13% vs 87.12%), and superior MSE on time-series forecasting tasks like Electricity (0.165 vs 0.175), demonstrating the effectiveness of our approach. The limitations and future directions are discussed in Appendix R.
@@
The authors have made extensive efforts to ensure the reproducibility of the empirical results reported in this paper. First, detailed dataset characteristics are documented in Appendix E.1. Second, comprehensive implementation details, covering network architectures, hyperparameter settings, training procedures, and evaluation metric explanations, are provided in Appendix LABEL:app:_implementation.
Acknowledgments
The authors would like to thank the anonymous reviewers for their valuable comments. This work was supported by the National Natural Science Foundation of China (No. 62076068).
@@ Appendix A Global Algorithm of BSD
28:   We calculate the loss of the last layer using cross-entropy: $\mathcal{L}_{\text{top}}\leftarrow\mathcal{L}_{\text{CE}}(\mathbf{v}_{L},\hat{\mathbf{v}}_{L})$
@@ Appendix B Additional Preliminaries
Spiking neurons can be categorized into two types based on their signal propagation mechanisms. One category, referred to as spike-based neurons, conveys information through spike trains. Discrete Spiking Neural Networks (SNNs) integrate spike inputs at each time step and generate a binary spike output when the integrated value exceeds a threshold. Continuous analog input signals are typically encoded (e.g., rate encoding, time encoding, or $\Delta$-encoding) into spike sequences. Representative models in this category include IF, PLIF(Fang et al., 2021), and LIF(Maass, 1997). Discrete SNNs are compatible with existing digital design processes, allowing the grouping and virtualization of large numbers of neurons on computational cores to achieve parallel architectures. These systems exhibit high scalability and efficient trainability, benefiting from sparse, event-driven computation with low power consumption during the inference phase. However, the discrete nature of spike signals limits their information capacity relative to traditional neuron models, which transmit continuous floating-point values directly(Pfeiffer and Pfeil, 2018).
The second type of spiking neuron architecture uses similar temporal equations as spiking neurons but propagates analog signals, referred to as analog-based neurons, such as LIAF(Wu et al., 2021). These signals can encode more information and exhibit lower response latency in certain tasks. However, hardware implementations of such neurons require analog circuits and because computations cannot be skipped during inference, the system exhibits higher power consumption (Hazan and Ezra Tsur, 2021).
@@ E.2.2 BSD-CNNs
Similarly to the MLP case, the feedback path, composed of Type 2 neurons, mirrors the structure of the feedforward path composed of Type 1 neurons. In the feedback path, we use upsampling as the reverse operation of the pooling used in the feedforward path, ensuring that the number of Type 1 and Type 2 neurons in feedforward and feedback paths remains equal across all layers.
@@ F.3 Ablation study on autoencoders
Appendix G Dynamics Of Synaptic Weight Alignment
@@
In this section, we examine the alignment dynamics between the feedforward synaptic weights ($\mathbf{W}$) and the backward synaptic weights ($\bm{\Theta}$) within a convolutional neural network trained using the BSD algorithm. The network was tasked with image classification on the CIFAR-10 dataset. We measured the cosine similarity between the weights connected to the basal dendrites of Type 1 and Type 2 neurons to quantify their alignment in each layer. In Figure 7, “enc1” through “enc5” denote the five feedforward convolutional layers and their corresponding backward upsampling layers, progressing from the shallowest to the deepest. The “output” label refers to the final fully connected layer. Figure 7 reveals that the weight alignment for the final fully connected layer undergoes a rapid initial increase, peaking within the first few epochs before stabilizing at a high cosine similarity value of approximately 0.7. In contrast, the weights of the convolutional and upsampling layers (“enc1” to “enc5”) consistently exhibit negligible alignment, with their cosine similarity values remaining close to zero throughout the entire training process. Such alignment behavior demonstrates that the feedforward and backward pathways do not converge toward symmetric weights under BSD. Thereby we confirm that the BSD algorithm adheres to criterion C1: Asymmetric synaptic weights for feedforward and feedback pathways.
Appendix H t-SNE Visualization of Learned Feature Representations in BSD-trained CNNs
@@
Appendix I Evaluation on Tiny-ImageNet
To evaluate the scalability of our proposed algorithm, we conducted experiments on the Tiny-ImageNet dataset. We compared our proposed BSD algorithm against both a standard backpropagation (BP) baseline and Dendritic Localized Learning (DLL). For a fair comparison, we used the same 5-layer CNN architecture across all three methods.
Table 8: Performance comparison on the Tiny-ImageNet dataset.
Method Accuracy (%)
BP 41.07
Dendritic Localized Learning (DLL) 17.10
BSD (Ours) 35.34
The results are summarized in Table 8. As shown, our BSD algorithm achieves an accuracy of 35.34%, substantially outperforming the other biologically plausible method, DLL.
Appendix J Ablation Study on Batch Size
To investigate the sensitivity of our BSD framework to the batch size, we conducted an ablation study using the BSD-trained CNN across the SVHN, CIFAR-10, and CIFAR-100 datasets. All other hyperparameters were kept consistent with the main experiments detailed in Appendix E.2. The results are summarized in Table 9.
Table 9: Impact of varying batch size on the performance of the BSD-trained CNN across multiple datasets.
Batch Size SVHN (%) CIFAR-10 (%) CIFAR-100 (%)
16 71.55 68.96 34.99
32 78.18 78.07 42.53
64 90.97 81.54 49.15
128 90.81 84.53 53.11
256 88.97 83.85 53.49
The results demonstrate a clear trend: the performance of our BSD framework generally improves as the batch size increases. At smaller batch sizes, the model’s performance is limited. This is an expected behavior for contrastive learning methods like ReCo, which require a sufficiently large and diverse set of negative samples within each batch to effectively shape the representation space. As the batch size increases, we observe substantial performance gains across all datasets, as larger batches provide a more robust estimate of the data distribution and a richer set of negative examples for the contrastive objective. We also observe that once the batch size reaches 128, the performance begins to stabilize. Further increasing the batch size to 256 yields minor fluctuations in accuracy. This suggests that a batch size of approximately 128 provides a sufficient number of negative samples for the model to achieve robust performance on these datasets. These findings confirm that our BSD framework scales favorably with batch size, which is consistent with the principles of contrastive learning.
Appendix K Energy Consumption Analysis
To quantify the energy efficiency of our trained models, we conducted a theoretical analysis of energy consumption during inference. The primary energy advantage of SNNs is realized in this phase, particularly on specialized neuromorphic hardware. Our BSD algorithm is a training methodology; the resulting model used for inference is a standard SNN. Therefore, a model trained with BSD fully retains the inherent energy-saving characteristics of SNNs.
We estimate the theoretical energy consumption per sample by following the methodology proposed by Yao et al. (2023b). For a given SNN layer $l$, the energy consumption is estimated as:
$\text{Energy}_{\text{SNN}}(l)=E_{\text{AC}}\times(T\times\gamma_{l}\times\text{FLOPs}(l)),$ (26)
where $T$ is the number of timesteps, $\gamma_{l}$ is the layer’s average firing rate, and $E_{\text{AC}}$ is the energy per accumulate operation. For an equivalent ANN layer $l$, the formula is:
$\text{Energy}_{\text{ANN}}(l)=E_{\text{MAC}}\times\text{FLOPs}(l),$ (27)
where $E_{\text{MAC}}$ is the energy per multiply-accumulate operation. The total energy for each model is the sum over all layers. We use the energy constants for a 45nm process from the literature ($E_{\text{AC}}=0.9$ pJ, $E_{\text{MAC}}=4.6$ pJ) for our analysis (Yao et al., 2023b). The analysis was performed on our BSD-trained 5-layer CNN and its equivalent ANN counterpart on the CIFAR-10 task. A layer-wise breakdown of the energy consumption for both models is provided in Table 10 and Table 11, with a final summary in Table 12.
Table 10: Layer-wise energy consumption analysis for the BSD-trained SNN during inference on a single sample from CIFAR-10.
Layer Name FLOPs Avg Firing Rate ($\gamma$) SOPs Energy (J)
Layer 1 (Conv2d) 3.54e+06 0.5006 7.09e+06 6.38e-06
Layer 2 (Conv2d) 3.77e+07 0.1519 2.29e+07 2.06e-05
Layer 3 (Conv2d) 1.89e+07 0.1815 1.37e+07 1.23e-05
Layer 4 (Conv2d) 9.44e+06 0.1943 7.34e+06 6.60e-06
Layer 5 (Conv2d) 4.72e+06 0.2037 3.85e+06 3.46e-06
Layer 6 (Linear) 5.12e+03 0.0803 1.65e+03 1.48e-09
Table 11: Layer-wise energy consumption analysis for the equivalent ANN-BP model during inference on a single sample from CIFAR-10.
Layer Name FLOPs Energy (J)
Layer 1 (Conv2d) 3.54e+06 1.63e-05
Layer 2 (Conv2d) 3.77e+07 1.74e-04
Layer 3 (Conv2d) 1.89e+07 8.68e-05
Layer 4 (Conv2d) 9.44e+06 4.34e-05
Layer 5 (Conv2d) 4.72e+06 2.17e-05
Layer 6 (Linear) 5.12e+03 2.36e-08
Table 12: Total energy consumption comparison per inference sample.
Model Total Operations per Sample Energy per Sample (J) Energy Reduction
ANN (BP) 7.43e+07 (FLOPs) 3.42e-04 -
SNN (BSD) 5.49e+07 (SOPs) 4.94e-05 85.5% $\downarrow$
This quantitative analysis confirms the substantial energy advantage of the SNN model. By leveraging sparse, event-driven computation, the SNN trained with our BSD method is approximately 85.5% more energy-efficient during inference than its architecturally identical ANN counterpart. This highlights that our biologically plausible training method yields models that are not only performant but also highly efficient for deployment.
Appendix L Robustness to Input Noise
To evaluate the robustness of our BSD framework against corrupted data, we conducted an experiment to assess the resilience of our algorithm to input noise. We used the model checkpoints pre-trained on the clean training sets of CIFAR-10 and SVHN. During the inference phase, we introduced corruption by adding random Gaussian noise (mean=0, std=0.05) to the input images. We then evaluated the performance of these clean-trained models on the noisy test data and compared it to a standard backpropagation (BP) baseline. The same 5-layer CNN architecture was used for both methods to ensure a fair comparison.
The results, summarized in Table 13, show the accuracy on both the original clean test set and the corrupted test set.
Table 13: Comparison of model accuracy on clean and noisy test data for CIFAR-10 and SVHN. Models were trained only on clean data. “Noisy Accuracy” corresponds to Accuracy Under Attack (AUA).
Dataset Method Clean Accuracy (%) Noisy Accuracy (%)
CIFAR-10 BP 87.18 70.89
BSD 83.67 67.06
SVHN BP 94.31 89.24
BSD 90.81 85.40
As expected, the performance of both methods degrades under noisy conditions. On both CIFAR-10 and SVHN, the drop in accuracy for BSD is comparable to that of BP. The results indicate that our approach learns robust and effective features.
We attribute this competitive robustness to the dual-objective nature of the BSD training process. The final task loss at the output layer pushes the network to extract discriminative information, while the layer-wise alignment objective simultaneously encourages the preservation of rich generative information from the input. This balance between two complementary goals acts as a form of regularization, guiding the model to learn more generalizable representations that are inherently more resilient to input perturbations, rather than learning brittle features that are only optimal for the clean data distribution.
Appendix M FFT Decomposition for Adaptive Loss
The FFT decomposition, used in our generation tasks, is a method to separate the low-frequency and high-frequency components of an image or a membrane voltage map, allowing us to apply an adaptive loss function. The process is as follows:
First, we apply a 2D Fast Fourier Transform to the spatial tensor (e.g., a batch of membrane voltage maps of shape $[B,C,H,W]$), converting it into its frequency-domain representation.
Second, to isolate these components, we construct frequency-domain masks. A low-pass filter is created by defining a circular region centered at the zero-frequency origin of the spectrum. Frequencies falling inside this radius are designated as low-frequency components, which represent the global structure and smooth areas of the image. Conversely, all frequencies outside this radius are designated as high-frequency components, which correspond to edges and fine-grained textures.
Finally, by element-wise multiplying the frequency-domain representation with these two masks, we separate it into two distinct tensors: one containing only low frequencies and another containing only high frequencies. This separation enables the adaptive loss computation mentioned in the main text. We calculate the ReCo loss independently on these two components but use a different penalty weight $\lambda$ for each. A smaller $\lambda$ is applied to the low-frequency components to maintain structural coherence, while a larger $\lambda$ is applied to the high-frequency components to more strongly enforce the preservation of sharp details and edge fidelity. This approach allows the model to better preserve fine details without sacrificing the image’s structural coherence.
Appendix N Computational Cost Analysis
To provide a comprehensive assessment of the computational costs associated with our proposed BSD algorithm, we conducted a comparative analysis against a standard Backpropagation (BP) baseline for SNNs. All experiments were performed on the CIFAR-10 dataset using a single NVIDIA GeForce RTX 2080 Ti GPU. We profiled memory consumption across different batch sizes to evaluate scalability. For both methods, we employed the 5-layer convolutional network architecture detailed in Appendix E.2 to ensure a fair comparison. The results for inference and training are presented separately below.
Inference Performance.
During inference for classification tasks, the BSD framework utilizes only its feedforward pathway (weights $\mathbf{W}$). Since this pathway is architecturally identical to a BP-trained SNN of the same configuration, the computational graph and parameters used during a forward pass are the same. As empirically confirmed in Table 14, this results in identical memory consumption across various batch sizes. This demonstrates that our biologically plausible training method introduces no computational overhead during the inference phase.
Table 14: Comparison of memory consumption during inference on CIFAR-10 across varying batch sizes.
Batch Size Method Inference Memory Consumption (MB)
32 BP 344
BSD 344
64 BP 944
BSD 944
128 BP 1826
BSD 1826
Training Performance.
During training, BSD must maintain two sets of weights ($\mathbf{W}$ and $\bm{\Theta}$) and store the activations for both the feedforward and backward pathways. Furthermore, the ReCo loss constructs a $B\times B$ affinity matrix, which introduces an additional memory usage component that scales quadratically with the batch size ($O(B^{2})$). However, the primary driver of BSD’s increased memory consumption is the necessity of storing the parameters and full activation maps for the second (backward) pathway, a characteristic inherent to our dual-network design. As detailed in Table 15, this architectural requirement leads to higher memory consumption compared to BP, which only stores one set of weights and the activations required for its backward pass.
Table 15: Comparison of memory consumption during training on CIFAR-10 across varying batch sizes.
Batch Size Method Training Memory Consumption (MB)
32 BP 950
BSD 1154
64 BP 1844
BSD 2852
128 BP 3324
BSD 5034
Time Complexity Analysis.
The empirical results are consistent with the theoretical time complexity of the algorithms. For a network with $L$ layers, $T$ timesteps, a batch size of $B$, and an average layer width of $D$, the complexities are as follows:
  • BPTT: The complexity is $O(L\cdot T\cdot B\cdot D^{2})$, derived from the forward pass and a symmetric backward pass through the unrolled computation graph.
  • BSD: The complexity is $O(L\cdot T\cdot B\cdot D^{2}+L\cdot B^{2}\cdot D)$. The first term, $O(L\cdot T\cdot B\cdot D^{2})$, accounts for propagation through the two pathways and is comparable to BPTT. The second term, $O(L\cdot B^{2}\cdot D)$, arises from computing the $B\times B$ affinity matrix for the layer-wise ReCo loss, making our training complexity more sensitive to batch size.
Appendix O Ablation Study on the Number of Timesteps
To more comprehensively investigate the influence of the number of timesteps ($T$) on our model’s performance, we expanded our ablation study across multiple datasets for both our CNN and RNN architectures. All other hyperparameters were kept consistent with those described in our main experiments. The results are summarized in Table 16 for image classification tasks and Table 17 for sequential regression tasks.
Table 16: Impact of varying timesteps ($T$) on the performance of BSD-trained CNNs across image classification datasets.
Timesteps ($T$) CIFAR-100 (%) SVHN (%) CIFAR-10 (%)
2 47.12 82.54 78.09
4 53.48 90.81 84.53
6 53.11 90.35 82.28
8 53.24 90.94 81.51
Table 17: Impact of varying timesteps ($T$) on the performance of BSD-trained RNNs across sequential regression tasks.
Timesteps ($T$) Harry Potter (acc $\uparrow$) Metr-la (mse $\downarrow$)
2 0.4185 0.1280
4 0.4169 0.1245
6 0.4168 0.1248
8 0.4212 0.1238
For both CNN and RNN models, performance generally improves when increasing the number of timesteps from $T=2$ to $T=4$. For timesteps greater than four ($T\geq 4$), performance tends to stabilize, exhibiting only minor fluctuations across most datasets and metrics. This suggests that four timesteps provide a robust and effective balance, allowing neurons to integrate sufficient information to form rich representations without introducing excessive temporal complexity. These expanded findings empirically justify our choice of $T=4$ as a robust and efficient setting across a variety of tasks and architectures.
Appendix P Ablation Study on Batch Normalization
To more broadly investigate the specific impact of batch normalization (BN), we expanded our ablation study to include the SVHN, CIFAR-10, and CIFAR-100 datasets. The experiments were conducted on our BSD-trained CNN, and the results are presented in Table 18.
Table 18: Impact of Batch Normalization (BN) on the performance of the BSD-trained CNN across multiple datasets.
Method SVHN (%) CIFAR-10 (%) CIFAR-100 (%)
BSD without BN 80.68 80.75 53.31
BSD with BN 90.81 84.53 53.48
The results show that incorporating BN improves performance across all tested datasets. This benefit is attributed to the ability of BN to stabilize the distribution of membrane potentials across layers. The normalization ensures that spike sparsity remains consistent, leading to more stable and effective learning.
Appendix Q Justification for the Choice of ReCo Loss
Here, we provide a more detailed justification for our choice of the Relaxed Contrastive (ReCo) loss over other contrastive alternatives like InfoNCE.
Our ablation study in Section 5.6 provides the empirical evidence, showing that BSD with ReCo loss consistently and significantly outperforms BSD-InfoNCE. This is particularly evident on complex datasets like CIFAR-100, where the performance margin is over 15 percentage points. The theoretical rationale for this superiority, especially for aligning the voltage signals in our BSD algorithm, is as follows.
The core objective at each hidden layer is to align the pre-spike membrane potential from the feedforward path ($\mathbf{v}_{i,k}$) with its corresponding supervisory voltage from the backward path ($\hat{\mathbf{v}}_{i,k}$). Simultaneously, the network must distinguish this pair from all negative pairs, where the voltage $\mathbf{v}_{i,k}$ is paired with supervisory signals from other samples in the batch ($\hat{\mathbf{v}}_{i,j}$ for $j\neq k$).
A standard InfoNCE loss enforces this distinction by applying a uniform repulsive force to all negative pairs. This compels the membrane voltage pattern generated for one sample to be strictly anti-correlated with the supervisory voltage patterns of all other samples. This is an overly restrictive constraint for high-dimensional membrane potentials, which can make the learning task unnecessarily difficult and potentially warp the feature space.
In contrast, the ReCo loss adopts a more targeted and flexible objective. It applies a penalty only when the voltage for one sample is confusingly similar to the supervisory signal of another (i.e., has a positive cosine similarity). If a negative pair’s voltage representations are already orthogonal or dissimilar, they incur no loss penalty. This allows the network to focus its learning capacity on separating the most confusable voltage signals, rather than expending effort pushing already distinct representations further apart. This fosters a richer and more flexible representational space for the membrane potentials, leading to the improved final performance that is empirically validated by our results.
Appendix R Limitations and Future Directions
R.1 Limitations